# EmphAssess : a Prosodic Benchmark on Assessing Emphasis Transfer in Speech-to-Speech Models

### v1

@@ Abstract
[…]
@@ 1 Introduction
Although scarce, there have been recent efforts made to establish benchmarks in the prosodic evaluation of speech models allowing models comparison, including evaluation corpora and pipelines, both at the global prosodic level (pragramtic information :  Lin et al. (2023) and at the local prosodic level (prosodic pauses: de Seyssel et al. (2023). Yet, there is a need for more benchmarks to cover other aspects of prosody, and all types of speech models.
In this work, we introduce the EmphAssess benchmark, which is focused on local prosody for speech-to-speech models and includes: (i) an new, automatic pipeline for emphasis evaluation that is modular, handles multiple languages and kinds of outputs (including paraphrases and translations, (ii) a novel dataset, the EmphAssess test set, for evaluating model emphasis preservation in English and Spanish according to our pipeline, and (iii) EmphaClass, an emphasis classifier that we finetuned with English data over an existing multilingual SSL model to support our pipeline.
@@ 4 The EmphAssess Dataset
The dataset generation started with a selection of transcripts from a list of handwritten transcripts with emphasis annotations previously created for company-internal Text-to-Speech purposes. Transcripts containing characters beyond letters or specific punctuation marks [^2] Retained punctuation characters include: [,:;.?!()] or those featuring proper nouns (identified using the NLTK toolkit; Bird 2006) were excluded. Moreover, we ensured a minimum of two distinct versions with different emphases for string identical sentences (those with matching word tokens but possibly differing emphasis position indices). This approach was adopted to mitigate any bias should a model exhibit a preference for emphasising a particular word over others. Finally, we filtered out transcripts that could face alignment challenges with emphasised words during translation. We set up an algorithm to assess the difficulty of aligning emphasised words in an English sentence with their counterparts in multiple target languages, using the SimAlign word-alignment tool Sabet et al. (2020). Simply put, if an emphasised word in the source matched consistently to a corresponding word across a list of other languages (German, French, Spanish, and Chinese), the sentence was labelled “easy”; otherwise, it was deemed “difficult.” Only “easy” transcripts were retained for our dataset. We were left with 913 distinct transcriptions (with varying emphases) derived from a pool of 299 unique transcriptions. We ensured that the distribution of transcripts was well balanced, in terms of where the emphasis was located.
@@
Finally, we compiled a dataset that is available as part of the benchmark. This dataset comprised four columns: an ‘id’ that denotes the unique identifier for each speech segment, a ‘src_sentence’ that contains the corresponding tokenised text transcript presented in list format, a ‘gold_emphasis’ that highlights the index of the emphasised word(s) also in list format, and a ‘voice’ that specifies the particular Expresso voice employed for the synthesis.
@@ 5 The EmphaAssess Evaluation Pipeline
The evaluation pipeline, as illustrated in Figure 1, is divided into two main stages. The first one (left panel) corresponds to the generation of utterances from the evaluated S2S model. That is, for each utterance from; the EmphAssess dataset, we need to generate the corresponding utterance output from the evaluated model. Hence, this inference stage is dependent on the model tested, and we will not expand on it here.
@@ 6.1 English models
For a topline evaluation, we matched the input utterances from EmphAssess with themselves (that is, we pretended the output utterances were the same as the input ones). This gave us an insight into the best achievable scores, with any potential loss in performance due to problems in the dataset or the various comparison stages. This topline produced an F1 score of 89%, indicating that our cascaded pipeline performs well. It should also be noted that we consider a baseline to yield scores of 0, corresponding to a model which does not encode emphasis and thus should not produce any emphasis.
@@ 6.2 Generalising the pipeline to S2S translation
As depicted in the right panel of Figure 3, the ‘topline,’ which aligns the English input with the synthesised Spanish voices as the output, achieved a score of 58%. While this result is reasonable, it notably lags behind the English topline. This decline may be attributed to various factors, including challenges in the synthesised voices, as we observed that our Spanish TTS voices do not emphasise as effectively as desired. Furthermore, issues in different stages of our automatic evaluation pipeline might contribute (for instance, the emphasis classifier’s performance is suboptimal in English, hinting that models trained in other languages might be more suitable). Additionally, linguistic differences could play a role, with Spanish emphasis potentially being less prominent than in English or conveyed through alternative means, possibly paraphrastically in the text itself. Nonetheless, having this topline facilitates the comparison of other models and the assessment of their relative performance. Subsequently, we evaluated the Seamless M4T model Barrault et al. (2023) in its English-to-Spanish translation capability, which yielded an F1 score of 14%. This result, akin to its English-to-English counterpart, suggests that the M4T model does not effectively capture emphasis.
@@ 6.3 Human Evaluation
To gauge human performance on the task, we conducted an evaluation with expert annotators. These annotators were presented with an utterance and its word-tokenised transcription, and were tasked with marking words they considered to be emphasised. Importantly, they were not obliged to mark any word as emphasised if they didn’t perceive any. This evaluation was carried out on a subset of the data, incorporating both English and Spanish utterances, with native annotators for each language. Figure 3 shows precision, recall, and F1 scores for English-to-English and English-to-Spanish, respectively. For English-to-Spanish, the human topline is set using the same Spanish utterances from the Spanish topline. These metrics were calculated by comparing the annotators’ identification of emphasis against the ‘gold standard’ annotation with which we synthesised the utterances.
Focusing first on the English dataset, the annotators achieved a commendable precision score of 86%, although this was offset by a lower recall score (50%). The lower recall could be attributed to annotators not perceiving emphasis in numerous sentences (Note: it is often harder to perceive emphasis in utterances taken out of context); nonetheless, the high precision score is encouraging. Turning our attention to the Spanish dataset, both recall and precision scores were lower. This aligns with our hypothesis that the quality of voice synthesis in Spanish was not up to par. It may also suggest that the nuances of emphasis might be linguistically specific, thereby differing between English and Spanish.
7 Discussion and Conclusion
@@
Additionally, we have open-sourced an emphasis-classification model that has been finetuned on English data. The model builds on a multilingual SSL architecture and has shown impressive accuracy in classifying emphasised speech in English on our dataset, along with reasonable performance in other languages (for further details, refer to the Appendix). The model’s robustness in English makes it a plausible starting point for finetuning classifiers in other languages, potentially minimising the volume of data needed for training. Interestingly, the fact that the successful results were achieved without retraining the encoder, suggests that the inherent features in the original XLS-R model were adequate for emphasis classification. Yet, it is important to underscore that we recommend using this classifier specifically for our defined use case (that is the presented benchmark and evaluation pipeline), as further testing is required to enhance its robustness and ensure its efficacy in detecting more nuanced forms of emphasis across other datasets.
@@
Lastly, our study is limited to binary categorisation of emphasis. Future endeavours could explore varying degrees of emphasis, although this would require more advanced models. For instance, capturing subtle differences in emphasis between the input and output of an S2S system could be a valuable addition to this line of research.
Beyond the limitations we have identified, there is an existing agenda for future research centring around the evaluation of prosody within SSL models. Firstly, on the subject of emphasis, we aim to scrutinise its functional role more closely—specifically, its ability to convey importance. We intend to investigate whether such a function is intrinsically represented within these models. Beyond emphasis, other aspects of prosody, such as turn-taking and speech grouping, merit attention. We are interested in determining whether these elements, too, are encoded within SSL models. Improved benchmarks and evaluations for these prosodic features could pave the way for the development of more expressive and nuanced models.

### v2

@@ Abstract
Maureen de Seysselthanks:   Currently at Apple  1,2    Antony D’Avirro1    Adina Williams1    Emmanuel Dupoux1,2 1Meta AI Research 2ENS, EHESS, CNRS, PSL University, France maureen.deseyssel@gmail.com       {adavirro,adinawilliams,dpx}@meta.com
@@ 1 Introduction
Although scarce, there have been recent efforts made to establish benchmarks in the prosodic evaluation of speech models allowing models comparison, including evaluation corpora and pipelines, both at the global prosodic level (pragmatic information :  Lin et al. (2023)) and at the local prosodic level (prosodic pauses: de Seyssel et al. (2023)). Yet, there is a need for more benchmarks to cover other aspects of prosody, and all types of speech models.
In this work, we introduce the EmphAssess benchmark, which is focused on local prosody for speech-to-speech models and includes: (i) a new, automatic pipeline for emphasis evaluation that is modular, handles multiple languages and kinds of outputs (including paraphrases and translations, (ii) a novel dataset, the EmphAssess test set, for evaluating model emphasis preservation in English and Spanish according to our pipeline, and (iii) EmphaClass, an emphasis classifier that we finetuned with English data over an existing multilingual SSL model to support our pipeline.
@@ 4 The EmphAssess Dataset
The dataset generation started with a selection of transcripts from a list of handwritten transcripts with emphasis annotations [^2] The emphasis could be applied to any sentence constituents, but it followed a contrastive pattern. previously created for company-internal Text-to-Speech purposes. Transcripts containing characters beyond letters or specific punctuation marks [^3] Retained punctuation characters include: [,:;.?!()] or those featuring proper nouns (identified using the NLTK toolkit; Bird 2006) were excluded, to ensure the translations are as straightforward as possible. Moreover, we ensured a minimum of two distinct versions with different emphases for string identical sentences (those with matching word tokens but possibly differing emphasis position indices). This approach was adopted to mitigate any bias should a model exhibit a preference for emphasising a particular word over others. Finally, we filtered out transcripts that could face alignment challenges with emphasised words during translation. We set up an algorithm to assess the difficulty of aligning emphasised words in an English sentence with their counterparts in multiple target languages, using the SimAlign word-alignment tool Sabet et al. (2020). Simply put, if an emphasised word in the source matched consistently to a corresponding word across a list of other languages (German, French, Spanish, and Chinese), the sentence was labelled “easy”; otherwise, it was deemed “difficult.” Only “easy” transcripts were retained for our dataset. We were left with 913 distinct transcriptions (with varying emphases) derived from a pool of 299 unique transcriptions. We ensured that the distribution of transcripts was well balanced, in terms of where the emphasis was located.
@@
Finally, we compiled a dataset that is available as part of the benchmark. This dataset comprised four columns: an id column that denotes the unique identifier for each speech segment, a src_sentence column that contains the corresponding tokenised text transcript presented in list format, a gold_emphasis column that highlights the index of the emphasised word(s) also in list format, and a voice column that specifies the particular Expresso voice employed for the synthesis.
@@ 5 The EmphaAssess Evaluation Pipeline
The evaluation pipeline, as illustrated in Figure 1, is divided into two main stages. The first one (left panel) corresponds to the generation of utterances from the evaluated S2S model. That is, for each utterance from the EmphAssess dataset, we need to generate the corresponding utterance output from the evaluated model. Hence, this inference stage is dependent on the model tested, and we will not expand on it here.
@@ 6.1 English S2S models
For a topline evaluation, we matched the input utterances from EmphAssess with themselves (that is, we pretended the output utterances were the same as the input ones). This gave us an insight into the best achievable scores, with any potential loss in performance due to problems in the dataset or the various comparison stages. This topline produced an F1 score of 89%, indicating that our cascaded pipeline performs well. It should also be noted that we consider chance-level to yield scores of 0, corresponding to a model which does not encode emphasis and thus should not produce any emphasis.
@@ 6.2 Generalising the pipeline to S2S translation
As depicted in the right panel of Figure 3, the ‘topline,’ which aligns the English input with the synthesised Spanish voices as the output, achieved a score of 58%. While this result is reasonable, it notably lags behind the English topline. This decline may be attributed to various factors, including challenges in the synthesised voices, as we observed that our Spanish TTS voices do not emphasise as effectively as desired. Furthermore, issues in different stages of our automatic evaluation pipeline might contribute (for instance, the Spanish emphasis classifier’s performance on spanish is not as optimal as its English counterpart on English data). Additionally, linguistic differences could play a role, with Spanish emphasis potentially being less prominent than in English or conveyed through alternative means, possibly paraphrastically in the text itself. Nonetheless, having this topline facilitates the comparison of other models and the assessment of their relative performance. Subsequently, we evaluated the Seamless M4T model Barrault et al. (2023) in its English-to-Spanish translation capability, which yielded an F1 score of 14%. This result, akin to its English-to-English counterpart, suggests that the M4T model does not effectively capture emphasis.
@@ 6.3 Human Evaluation
To gauge human performance on the task, we conducted an evaluation with expert annotators. These annotators were presented with an utterance and its word-tokenised transcription, and were tasked with marking words they considered to be emphasised. Importantly, they were not obliged to mark any word as emphasised if they didn’t perceive any. This evaluation was carried out on a subset of the data, incorporating both English and Spanish utterances, with native annotators for each language. Figure 3 shows precision, recall, and F1 scores for English-to-English and English-to-Spanish, respectively [^5] For English-to-Spanish, the human topline is set using a subset of the Spanish utterances synthesized the Spanish topline. These metrics were calculated by comparing the annotators’ identification of emphasis against the ‘gold standard’ annotation with which we synthesised the utterances.
Focusing first on the English dataset, the annotators achieved a commendable precision score of 86%, although this was offset by a lower recall score (50%). The lower recall could be attributed to annotators not perceiving emphasis in numerous sentences (Note: it is often harder to perceive emphasis in utterances taken out of their general, wider context); nonetheless, the high precision score is encouraging. Turning our attention to the Spanish dataset, both recall and precision scores were lower. This aligns with our hypothesis that the quality of voice synthesis in Spanish was not up to par - with the larger drop of recall compared to the topline could be explained by the Spanish emphasis classifier model picking up very subtle cues that are not obvious to the human ear. It may also suggest that the nuances of emphasis might be linguistically specific, thereby differing between English and Spanish.
7 Conclusion
@@
Additionally, we have open-sourced an emphasis-classification model that has been finetuned on English data. The model builds on a multilingual SSL architecture and has shown impressive accuracy in classifying emphasised speech in English on our dataset, along with reasonable performance in other languages (for further details, refer to the Appendix). The model’s robustness in English makes it a plausible starting point for finetuning classifiers in other languages, potentially minimising the volume of data needed for training. Interestingly, the fact that the successful results were achieved without retraining the encoder, suggests that the inherent features in the original XLS-R model were adequate for emphasis classification.
There is an existing agenda for future research centring around the evaluation of prosody within SSL models. Firstly, on the subject of emphasis, we aim to scrutinise its functional role more closely—specifically, its ability to convey importance. We intend to investigate whether such a function is intrinsically represented within these models. Beyond emphasis, other aspects of prosody, such as turn-taking and speech grouping, merit attention. We are interested in determining whether these elements, too, are encoded within SSL models. Improved benchmarks and evaluations for these prosodic features could pave the way for the development of more expressive and nuanced models.
8 Limitations
While pioneering in its approach to evaluating emphasis in S2S models, our study encounters certain limitations. First, the emphasis classifier presented in this paper was made to be used with this exact dataset, and we recommend constraining its use to this particular use case (that is, with the presented benchmark and evaluation pipeline). Indeed, further testing is required to enhance its robustness and ensure its efficacy in detecting more nuanced forms of emphasis across other datasets.
@@ 8 Limitations
Another limitation of our work lies in the use of synthesised speech to create our dataset. While this approach provides a more controlled and consistent dataset—for instance, by enabling the synthesis of identical textual content with varying word emphases and voices—it may fail to capture the full range of characteristics found in natural speech. Consequently, this limitation could affect how well the benchmark results can be applied to practical use cases.
Lastly, our study is currently limited to binary categorisation of emphasis. Future endeavours could explore varying degrees of emphasis, although this would require more advanced models. For instance, capturing subtle differences in emphasis between the input and output of an S2S system could be a valuable addition to this line of research.